What is the critical angle for total internal reflection at the interface between
diamond and air?
1
Expert's answer
2020-11-16T07:52:20-0500

The critical angle for total internal reflection is given by:


"\theta_c = \arcsin\left( \dfrac{n_2}{n_1} \right)"

where "n_2 = 1" is the refractive index for air and "n_1\approx 2.42" is the refractive index of diamond. Thus, obtain:


"\theta_c = \arcsin\left( \dfrac{1}{2.42} \right) \approx 24.4\degree"

Answer. "24.4\degree".